### v3

# Symmetries of Hypergraphs and Some Invariant Subspaces of Matrices Associated with Hypergraphs

Abstract.
Here, the structural symmetries of a hypergraph are represented through equivalence relations on the vertex set of the hypergraph. A matrix associated with the hypergraph may not reflect a specific structural symmetry. In the context of a given symmetry within a hypergraph, we investigate a collection of matrices that encapsulate information about the symmetry. Our investigation reveals that certain structural symmetries in a hypergraph manifest observable effects on the eigenvalues and eigenvectors of designated matrices associated with the hypergraph. We identify invariant subspaces of these specific matrices as a consequence of symmetries present in the hypergraph. These invariant subspaces elucidate analogous behaviours observed in certain clusters of vertices during random walks and other dynamical processes on the hypergraph.
Key words and phrases:
Hypergraphs; Hypergraphs and linear algebra; Symmetries in hypergraphs; Invariant subspaces of hypergraph matrices; Eigenvalues and eigenvectors of hypergraph matrices.
2020 Mathematics Subject Classification:
Primary 05C65 05C50, 05C69, 05C70 ; Secondary 05C75,05C07,05C30
1. Introduction
A hypergraph $H$ is an ordered pair $(V(H),E(H))$ such that $V(H)$ is a non-empty set and $E(H)$ is a family of non-empty subsets of $V(H)$. Each element of $V(H)$ is called a vertex in $H$, and $V(H)$ is called the vertex set of $H$. Any element of $E(H)$ is called a hyperedge, and $E(H)$ is called a hyperedge set in $H$. A hyperedge $e\in E(H)$ is called a loop if $|e|=1$. If $V(H)$ is a finite set, $H$ is called a finite hypergraph. Here, we only consider finite hypergraphs without loops. For any $i\in V(H)$, we define the star of the vertex $i$ as $E_{i}(H)=\{e\in E(H):i\in e\}$. We denote the collection of stars for all vertices in $H$ as $\mathfrak{E}(H)$. That is, $\mathfrak{E}(H)=\{E_{i}(H):i\in V(H)\}$.
A hypergraph $H$ is called $m$-uniform if $|e|=m$ for all $e\in E(H)$. Thus, graphs are $2$-uniform hypergraphs. The relationship between the structural symmetry of graphs and the spectra of matrices associated with graphs has been well-studied. These symmetries can be represented by graph homomorphism, graph covering, quotient graph, twin vertices, graph automorphisms, and quotient matrices. Traces of these symmetries in the spectra of matrices related to the graph are explored in [6, 10, 13, 11, 14, 9]. Various matrices, such as adjacency matrix, Laplacian matrix, signless Laplacian matrix etc. are used to investigate hypergraphs ([1, 16, 3, 12, 15, 7]). Given a structural symmetry of a hypergraph, a natural question is how to represent the symmetry. For example, in Figure 1, any pair of vertices in $U=\{5,6,7\}\subset V(H)$ have the same stars. Thus, $U$ is a symmetric substructure in $H$. In [2], we represent this type of symmetry using equivalence relations on the set of vertices of the hypergraph. Given any equivalence relation $\mathfrak{R}$ on the vertex set, all the vertices belonging to an $\mathfrak{R}$-equivalence class are similar in some sense. Thus, $\mathfrak{R}$ represents a symmetry of the hypergraph. Since any matrix associated with a hypergraph $H$ may contain only partial information about $H$, given a particular equivalence relation $\mathfrak{R}$ on the vertex set $V(H)$, a matrix related to $H$ may or may not encode the information about the symmetry represented by $\mathfrak{R}$. Thus, it is natural to ask what family of matrices associated with a hypergraph contains the information about the equivalence relation. In [3], we introduced some general operators related to hypergraphs. In [2], for any equivalence relation on $V(H)$, we have introduced the notion of equivalence relation compatible operators, a family of operators that can encrypt the traces of symmetry represented by the equivalence relation. We show some particular symmetries of hypergraphs, which are captured by the general operators associated with the hypergraph. In this article, we find some larger classes of matrices that can encrypt information about these symmetries in terms of their invariant subspaces and as well as their eigenvalues. We show that these families of matrices contain the adjacencies and Laplacian matrices associated with hypergraphs studied so far ( see 2.8 and 2.9). In this study, we demonstrate that certain symmetries inherent in hypergraphs give rise to invariant subspaces of matrices associated with the hypergraph. For example, here we find star-dependent matrices that encode the effect of the symmetry represented by the equivalence relation $\mathfrak{R}_{s}(H)$ that depends on the star of the vertices. We also have the edge-sum matrices related to the equivalence relation $\mathfrak{R}_{2}(H,M_{H})$. The effect of these symmetries, represented by the equivalence relation $\mathfrak{R}_{s}(H)$ and $\mathfrak{R}_{2}(H,M_{H})$ can be traced in the eigenvalues and eigenvectors of the star-dependent matrices and edge-sum matrices, respectively. We show that these effects on the eigenvectors provide invariant subspaces of related matrices. These invariant subspaces cause similar behaviors in clusters of vertices during dynamical processes in hypergraphs.
The Section 2 starts with an equivalence relation $\mathfrak{R}_{M_{H}}$ on the vertex set $V(H)$ of a hypergraph $H$. Let $\mathbb{R}^{V(H)}$ be the set of all the functions $x:V(H)\to\mathbb{R}$. For any $x\in\mathbb{R}^{V(H)}$, $\psi_{ij}(x)\in\mathbb{R}^{V(H)}$ is such that $(\psi_{ij}(x))(i)=x(j)$, $(\psi_{ij}(x))(j)=x(i)$, and $(\psi_{ij}(x))(k)=x(k)$ for all $k\in V(H)\setminus\{i,j\}$. Given a square matrix $M_{H}$ of order $|V(H)|$, two vertices $i,j\in V(H)$ are $\mathfrak{R}_{M_{H}}$-related if $M_{H}(\psi_{ij}(x))=\psi_{ij}(M_{H}x)$ for any $x\in\mathbb{R}^{V(H)}$. We find the symmetries encoded by the equivalence relation $\mathfrak{R}_{M_{H}}$ leave traces in the eigenvalues, eigenvectors, and the invariant subspaces of $M_{H}$ (see 2.2, 2.3). The equivalence relation $\mathfrak{R}_{M_{H}}$ is related to the matrix $M_{H}$. Using 2.4 and $\mathfrak{R}_{M_{H}}$, we find that the structural symmetries of a hypergraph encoded by the equivalence relations $\mathfrak{R}_{s}(H)$, and $\mathfrak{R}_{2}(H,M_{H})$ also leave traces in the eigenvalues, eigenvectors, and invariant subspaces of some family of matrices, namely star-dependent matrices, edge-sum matrices, edge-cardinality matrices. Two vertices $i,j\in V(H)$ are $\mathfrak{R}_{s}(H)$-related if their stars are equal. Each $\mathfrak{R}_{s}(H)$-equivalence class is called a unit in the hypergraph $H$. A star dependent matrix (defined in 2.7) is a matrix of order $|V(H)|$, whose $i,j$-th element depends only on the star of $i$, and $j$. We show that if $M_{H}$ is a star-dependent matrix, then the structural symmetries in the units lead us to invariant subspaces of $M_{H}$ (see 2.10, 2.11,2.12). An edge-sum matrix is a star-dependent matrix of order $|V(H)|$ such that for all $i,j\in V(H)$, the $i,j$-th entry of the matrix is a weighted sum on the intersection of stars, $E_{i}(H)\cap E_{j}(H)$. For an edge-sum matrix $M_{H}=\left(m_{ij}\right)_{i,j\in V(H)}$, if two vertices $i,j\in V(H)$ are $\mathfrak{R}_{2}(H,M_{H})$-related, then due to a structural symmetry, represented by the bijections between the stars $E_{i}(H)$, $E_{j}(H)$ we have $m_{ik}=m_{jk}$, and $m_{ki}=m_{kj}$. We find invariant subspaces of edge-sum matrices due to the symmetry associated with $\mathfrak{R}_{2}(H,M_{H})$ (see 2.14, and 2.15). We use our results to explain similar behaviours in the clusters of vertices of hypergraphs in random walks and other dynamical processes on hypergraphs.
2. Invariant Subspaces Due to Transpositions
Let $H$ be a hypergraph and $M_{H}$ be a square matrix whose rows and columns are indexed by the set of vertices $V(H)$ in $H$, that is $M_{H}=\left(m_{ij}\right)_{i,j\in V(H)}$. For any square matrix $M$, we denote the set of all eigenvalues of $M$ as $\sigma_{M}$. Let $V_{\lambda}(M)$ be the eigenspace of $\lambda\in\sigma_{M}$, that is, $V_{\lambda}(M)$ is the vector space generated by $\{x:Mx=\lambda x\}$. For two distinct $i,j\in V(H)$, the transposition $\phi_{ij}:V(H)\to V(H)$ is defined as
[…]
The bijection $\phi_{ij}$ leads us to the map $\psi_{ij}:\mathbb{R}^{V(H)}\to\mathbb{R}^{V(H)}$ defined by $\psi_{ij}(x)=x\circ\phi_{ij}$. If two square matrices, $A$ and $B$ commutes, the eigenspace of any eigenvalue of $A$ is invariant under the action of $B$. As an example, any eigenspace of the adjacency matrix of a graph $G$ is invariant under the action of any automorphism of $G$ ( [8, Section-1]). This observation motivates us to consider the following equivalence relation.
$\mathfrak{R}_{M_{H}}:=\{(i,j)\in V(H)\times V(H):M_{H}(\psi_{ij}(x))=\psi_{ij}(M_{H}x)\text{~{}for~{}all}~{}x\in{\mathbb{R}}^{V(H)}\}.$
Any positive real-valued function $f:V(H)\to(0,\infty)$ provides an inner product $(\cdot,\cdot)_{f}$ on $\mathbb{R}^{V(H)}$, defined as
$(x,y)_{f}=\sum\limits_{i\in V(H)}f(i)x(i)y(i)\text{~{}for all~{}}x,y\in\mathbb{R}^{V(H)}.$
We say $M_{H}$ is an $f$-symmetric matrix if $(M_{H}x,y)_{f}=(x,M_{H}y)_{f}$ for all $x,y\in\mathbb{R}^{V(H)}$. For any $x\in\mathbb{R}^{V(H)}$, the support of $x$ is $supp(x)=\{i\in V(H):x(i)\neq 0\}$. We define the following for any $U\subseteq V(H)$.
(1)$S_{U}=\{x\in\mathbb{R}^{V(H)}:supp(x)\subseteq U,\sum\limits_{i\in U}x(i)=0\}$.
(2)For any $f:V(H)\to(0,\infty)$, $C^{f}_{U}=\{x\in\mathbb{R}^{V(H)}:f(i)x(i)=c_{x},\text{~{}a real constant for all~{}}i\in U\}$. If $f(i)=1$ for all $i\in V(H)$, then we denote $C^{f}_{U}$ as $C_{U}=\{x\in\mathbb{R}^{V(H)}:x(i)=c_{x}\text{~{}for all~{}}i\in U\}$.
(3)The characteristic function $\chi_{{}_{U}}:\mathbb{R}^{V(H)}\to\mathbb{R}^{V(H)}$ of $U$ is defined as $\chi_{{}_{U}}(i)=\begin{cases}1&\text{~{}if~{}}i\in U,\\ 0&\text{otherwise.}\end{cases}$ For two distinct $i,j\in V(H)$, the function $x_{ij}=\chi_{\{j\}}-\chi_{\{i\}}$ is referred to as a Faria vector (see [4, p.46, Section 3.7]). For any $U=\{i_{0},i_{1},\ldots,i_{n-1}\}\subseteq V(H)$, the collection $\{x_{i_{j}i_{0}}:j=1,\ldots,n-1\}$ forms a basis of $S_{U}$. Therefore, $S_{U}$ is a subspace of dimension $|U|-1$. The next result shows a relation between $S_{U}$ and $C^{f}_{U}$. For any subspace $\mathcal{V}$ of $\mathbb{R}^{V(H)}$, we denote the orthogonal complement of $\mathcal{V}$ with respect to the inner product $(\cdot,\cdot)_{f}$ is $\mathcal{V}^{(\perp,f)}$.
Proposition 2.1.
Let $H$ be a hypergraph. For any $U\subseteq V(H)$ with $|U|>1$, and any function $f:V(H)\to(0,\infty)$, the orthogonal complement of $S_{U}$ with respect to the inner product $(\cdot,\cdot)_{f}$, $S_{U}^{(\perp,f)}=C^{f}_{U}$.
Proof.
Let $U=\{i_{0},i_{1},\ldots,i_{n-1}\}\subseteq V(H)$. For any $y\in S_{U}^{(\perp,f)}$, if $i,j(\neq i)\in U$, then
$\sum_{u\in V(H)}f(u)y(u)x_{ij}(u)=0,$
and so $f(i)y(i)=f(j)y(j)$ for all $i,j(\neq i)\in U$. Therefore, $y\in C^{f}_{U}$, and thus, $S_{U}^{(\perp,f)}\subseteq C^{f}_{U}$. Conversely, for any $y\in C_{U}^{f}$, $f(i)y(i)=c_{y}$ for all $i\in U$. Thus, if $x\in S_{U}$, then
$\sum\limits_{i\in V(H)}f(i)x(i)y(i)=\sum\limits_{i\in U}f(i)x(i)y(i)=c_{y}\sum\limits_{i\in U}x(i)=0,$
and therefore, $y\in S_{U}^{(\perp,f)}$. Thus, $C^{f}_{U}\subseteq S_{U}^{(\perp,f)}$ and the result follows. ∎
For $(i,j)\in\mathfrak{R}_{M_{H}}$ a symmetry between $i$ and $j$ in $H$ is encoded in $M_{H}$ in such a way that the action of $M_{H}$ commutes with the transposition between the vertices $i$ and $j$. In the next result, we show that this symmetry contributes to the spectra of $M_{H}$.
Proposition 2.2.
Let $H$ be a hypergraph. For any $\mathfrak{R}_{M_{H}}$-equivalence class $W$ with $|W|\geq 2$, there exists $\lambda_{W}\in\sigma_{M_{H}}$, such that $S_{W}$ is a subspace of $V_{\lambda_{W}}(M_{H})$.
Proof.
Suppose that $W=\{i_{0},i_{1},\ldots,i_{n-1}\}$. For any $i,j\in W$ we have $\phi_{ij}(x_{ij})=-x_{ij}$. Thus, if $u\in V(H)\setminus\{i,j\}$ then
$(\phi_{ij}(M_{H}x_{ij}))(u)=M_{H}(\phi_{ij}(x_{ij}))(u)$
$\implies(M_{H}x_{ij})(u)=0,$
and $(\phi_{ij}(M_{H}x_{ij}))(i)=(M_{H}(\phi_{ij}(x_{ij})))(i)$ implies $(M_{H}x_{ij})(j)=-(M_{H}x_{ij})(i)$. Therefore, $M_{H}x_{ij}=-[(M_{H}x_{ij})(i)]x_{ij}$. For three distinct $i,j,k\in W$, since $(M_{H}(\phi_{jk}(x_{ij})))(i)=(\phi_{jk}(M_{H}(x_{ij})))(i)$, we have $(M_{H}x_{ik})(i)=(M_{H}(x_{ij}))(i)$. Therefore, $[(M_{H}x_{i_{0}i_{j}})(i_{0})]=\lambda_{W}$, a constant for all $j=1,\ldots,n-1$ and,
$M_{H}x_{i_{0}i_{j}}=-\lambda_{W}x_{i_{0}i_{j}}.$ (2.1)
Since $\{x_{i_{0}i_{j}}:j=1,\ldots,n-1\}$ is a basis for $S_{W}$, we have $S_{W}$ is a subspace $V_{\lambda_{W}}(M_{H})$. This completes the proof. ∎
Let $M_{H}$ be a matrix associated with a hypergraph $H$. A subspace $W$ of the vector space $\mathbb{R}^{V(H)}$ is called an invariant subspace of $M_{H}$ if $M_{H}y\in W$ for all $y\in W$. In the following Theorem, for any matrix $M_{H}$, each $\mathfrak{R}_{M_{H}}$-equivalence class with cardinality at least $2$ corresponds to an invariant subspace of $M_{H}$.
Theorem 2.3.
Let $H$ be a hypergraph, $f:V(H)\to(0,\infty)$ be a function, and $M_{H}$ be an $f$-symmetric matrix. For any $\mathfrak{R}_{M_{H}}$-equivalence class $W$ with $|W|\geq 2$, the vector space $C_{W}^{f}$ is an invariant subspace of $M_{H}$.
Proof.
By 2.2, for an $\mathfrak{R}_{M_{H}}$-equivalence class $W$ with $|W|\geq 2$, there exists an eigenvalue $\lambda_{W}$ of $M_{H}$, such that, $S_{W}$ is a subspace of the eigenspace $V_{\lambda}(M_{H})$. Thus, $S_{W}$ is an invariant subspace of $M_{H}$, and since $C_{W}^{f}=S_{W}^{(\perp,f)}$ by 2.1, the vector space $C_{W}^{f}$ is an invariant subspace of $M_{H}$. ∎
2.3 indicates that for any hypergraph $H$ and any matrix $M_{H}$ indexed by $V(H)$, if $\{x_{n}:V(H)\to\mathbb{R}:n\in\mathbb{N}\}$ is governed by the difference equation
$x_{n+1}=M_{H}x_{n},$ (2.2)
and there exists $k\in\mathbb{N}$ such that $x_{k}$ is constant in a $\mathfrak{R}_{M_{H}}$-equivalence class $W$ then so is $x_{n}$ for all $n\geq k$. Since many phenomena on hypergraph can be described using difference equation similar to the Equation 2.2, in the remaining part of this article, we use 2.3 to explore the effect of symmetries of hypergraph on these phenomena.
The equivalence relation $\mathfrak{R}_{M_{H}}$ is related to the matrix $M_{H}$. Now, we relate some structural symmetries of hypergraphs with $\mathfrak{R}_{M_{H}}$. Let $\mathfrak{R}_{1}$ and $\mathfrak{R}_{2}$ denote two equivalence relations on the vertex set $V(H)$. The relation $\mathfrak{R}_{1}$ is considered finer than $\mathfrak{R}_{2}$ if $\mathfrak{R}_{1}\subseteq\mathfrak{R}_{2}$. This implies that if $(i,j)\in\mathfrak{R}_{1}$, then $(i,j)\in\mathfrak{R}_{2}$, indicating that for any $\mathfrak{R}_{1}$-equivalence class $W_{1}$, there exists a corresponding $\mathfrak{R}_{2}$-equivalence class $W_{2}$ such that $W_{1}\subseteq W_{2}$.
As demonstrated in 2.3, it is established that $C^{f}_{W}$ is an invariant subspace of $M_{H}$, where $W$ represents an $\mathfrak{R}_{M_{H}}$-equivalence class. We now extend this result to assert that the same holds for any $\mathfrak{R}$-equivalence class $W$, where $\mathfrak{R}$ is an equivalence relation finer than $\mathfrak{R}_{M_{H}}$.
Theorem 2.4.
Let $H$ be a hypergraph, $f:V(H)\to(0,\infty)$ be a function, and $M_{H}$ be an $f$-symmetric matrix. If $\mathfrak{R}$, an equivalence relation on $V(H)$, is finer than $\mathfrak{R}_{M_{H}}$ then for any $\mathfrak{R}$-equivalence class $W$ with cardinality $|W|\geq 2$, the vector space $C_{W}^{f}$ is an invariant subspace of $M_{H}$.
Proof.
Since for any $\mathfrak{R}$-equivalence class $W$, there exists an $\mathfrak{R}_{M_{H}}$-equivalence class $W^{\prime}$ such that $W\subseteq W^{\prime}$, and thus, $S_{W}$ is a subspace of $S_{W^{\prime}}$. So, Equation 2.1 concludes that $S_{W}$ is a subspace generated by some eigenvectors of the eigenvalue $\lambda_{W^{\prime}}\in\sigma_{M_{H}}$. Thus, $S_{W}$ is an invariant subspace of $M_{H}$, and so is $C_{W}^{f}$. ∎
Now, we introduce some equivalence relations related to structural symmetries of $H$, which are finer than $\mathfrak{R}_{M_{H}}$. We use 2.4 to explore the relation of these structural symmetries of the hypergraph $H$ with the matrix $M_{H}$.
2.1. Units, Star Dependent Matrices and their Invariant Subspaces
Units in a Hypergraph
In a hypergraph $H$, units are the maximal subsets of vertices having the same star.
Definition 2.5 (Units, [2]).
Let $H$ be the hypergraph and $\mathfrak{R}_{s}(H)$ be an equivalence relation on $V(H)$ defined by
$\mathfrak{R}_{s}(H)=\{(i,j)\in V(H)\times V(H):E_{i}(H)=E_{j}(H)\}.$
Then, each $\mathfrak{R}_{s}(H)$-equivalence class in $H$ is called a unit.
A unit in $H$ corresponds to an $E\subseteq E(H)$, such that the unit is $\{i\in V(H):E_{i}(H)=E\}$, the maximum collection of vertices $i$, with $E_{i}(H)=E$. We denote the unit as $W_{E}$ and refer to $E$ as the generating set of the unit $W_{E}$. We denote the collection of all the units in a hypergraph $H$ as $\mathfrak{U}(H)$.
{tikzpicture}
Figure 1. A hypergraph $H$ wherein all the units of $H$ are illustrated by shaded with dots.
Example 2.6.
Consider the hypergraph $H$ in Figure 1, with $V(H)=\{i\in\mathbb{N}:1\leq i\leq 18\}$, and $E(H)=\{c,d,e,f,g,h\}$, where $c=\{8,13,14,15,16\}$, $d=\{3,4,8\}$, $e=\{1,2,8\}$, $f=\{5,6,7,8\}$, $g=\{8,9,10,11,12\}$, $h=\{8,12,13,14,17,18\}$. In $H$, the units are $W_{E_{1}}=\{1,2\}$, $W_{E_{2}}=\{5,6,7\}$, $W_{E_{3}}=\{9,10,11\}$, $W_{E_{4}}=\{12\}$, $W_{E_{5}}=\{13,14\}$, $W_{E_{6}}=\{15,16\}$, $W_{E_{7}}=\{17,18\}$, $W_{E_{8}}=\{8\}$, and $W_{E_{9}}=\{3,4\}$. Here the generating sets are $E_{1}=\{e\}$, $E_{2}=\{f\}$, $E_{3}=\{g\}$, $E_{4}=\{g,h\}$, $E_{5}=\{c,h\}$, $E_{6}=\{c\}$, $E_{7}=\{h\}$, and $E_{8}=E(H)$.
Now, we present a set of matrices associated with a hypergraph $H$ that captures the effects of symmetry within the units of $H$.
Definition 2.7 (Star dependent matrix).
Let $H$ be a hypergraph. We call a matrix $M_{H}=\left(M_{ij}\right)_{i,j\in V(H)}$ a star dependent matrix if there exists two functions $g:\mathfrak{E}(H)\times\mathfrak{E}(H)\to\mathbb{R}$, and $h:\mathfrak{E}(H)\to\mathbb{R}$ such that,
[…]
We refer to the functions $f$ and $g$ as the diagonal and non-diagonal star-function of the star-dependent matrix $M_{H}$, respectively.
Example 2.8 (Adjacency Matrix).
  1. (1)
    According to [5], the adjacency matrix of a hypergraph $H$ is a square matrix $A^{(1)}$, whose rows and columns are indexed by $V(H)$ and for $i,j(\neq i)\in V(H)$, the $(i,j)$-th entry is $A^{(1)}_{ij}=|E_{i}(H)\cap E_{j}(H)|$, the co-degree of $i$ and $j$, and the diagonal position is $A^{(1)}_{ii}=0$ for all $i\in V(H)$. Since $A^{(1)}_{ij}=|E_{i}(H)\cap E_{j}(H)|$ depends only on $E_{i}(H)$, $E_{j}(H)$ and $A^{(1)}_{ii}=0$ is a constant, $A^{(1)}$ is a star-dependent matrix.
  2. (2)
    Similarly, the adjacency matrix $A^{(2)}=\left(A^{(2)}_{ij}\right)_{i,j\in V(H)}$ of $H$, introduced in [1], defined by
[…]
    is also a star-dependent matrix.
  3. (3)
    The normalized adjacency matrix $\mathcal{A}=\left(\mathcal{A}_{ij}\right)_{i,j\in V(H)}$ of $H$, defined by
[…]
    is a star-dependent matrix.
Example 2.9 (Laplacian Matrix).
  1. (1)
    The Laplacian matrix $L^{(1)}=D^{(1)}-A^{(1)}$ is a star-dependent matrix, where $D^{(1)}$ is the diagonal matrix of order $|V(H)|$, defined by $D^{(1)}_{ii}=(A^{(1)}\chi_{V(H)})(i)$. This variation of hypergraph Laplacian $L^{(1)}$ is used in [12].
  2. (2)
    Another variation of Laplacian $L^{(2)}=D^{(2)}-A^{(2)}$, introduced in [1], is also a star-dependent matrix.
  3. (3)
    The normalized Laplacian $\Delta=I_{|V(H)|}-A^{(2)}$ introduced in [1], is a star-dependent matrix. Here $I_{|V(H)|}$ is the identity matrix of order $|V(H)|$.
Theorem 2.10.
If $H$ is a hypergraph and $M_{H}$ is a star-dependent matrix, then $\mathfrak{R}_{s}(H)$ is finer than $\mathfrak{R}_{M_{H}}$.
Proof.
For $i,j\in V(H)$, assuming $(i,j)\in\mathfrak{R}_{s}(H)$, it is enough to show $(i,j)\in\mathfrak{R}_{M_{H}}$. Since $(i,j)\in\mathfrak{R}_{s}(H)$ we have $E_{i}(H)=E_{j}(H)$. Since $M_{H}=\left(m_{ij}\right)_{ij\in V(H)}$ is a star-dependent matrix, $E_{i}(H)=E_{j}(H)$ leads us to $m_{ki}=m_{kj}$, and $m_{ik}=m_{jk}$ for all $k\in V(H)$. Thus,
$(\psi_{ij}(M_{H}x))(r)$ =$=\begin{cases}\sum\limits_{k\in V(H)}m_{rk}x(k)&\text{~{}if~{}}r\notin\{i,j\}\\ \sum\limits_{k\in V(H)}m_{ik}x(k)&\text{~{}if~{}}r=j\\ \sum\limits_{k\in V(H)}m_{jk}x(k)&\text{~{}if~{}}r=i.\end{cases}$
$=(M_{H}x)(r),$
and
$(M_{H}(\psi_{ij}(x)))(r)=\sum\limits_{k\in V(H)\setminus\{i,j\}}m_{rk}x(k)+m_{ri}x(j)+m_{rj}x(i)=(M_{H}x)(r).$
Therefore, $\psi_{ij}(M_{H}x)=M_{H}(\psi_{ij}(x))$, and $(i,j)\in\mathfrak{R}_{M_{H}}$. ∎
For any star-dependent matrix $M_{H}$, 2.2 and the proof of 2.10 show that if $W_{E}$ is a unit in $H$, then for all $i,j(\neq i)\in W_{E}$, $\lambda_{W_{E}}=M_{H}x_{ij}(i)$ is an eigenvalue of $M_{H}$ and $x_{ij}\in V_{\lambda_{W_{E}}}(M_{H})$. Therefore, $S_{W_{E}}$ is a subspace of $V_{\lambda_{W_{E}}}(M_{H})$. Thus, 2.4 leads us to the following result on an invariant subspace of star-dependent matrices.
Corollary 2.11 (Unit-invariant subspace).
Let $H$ be a hypergraph, $f:V(H)\to(0,\infty)$ be a function, and $M_{H}$ be an $f$-symmetric, star-dependent matrix indexed by $V(H)$. For any $W\in\mathfrak{U}(H)$ with $|W|\geq 2$, the vector space $C_{W}^{f}$ is an invariant subspace of $M_{H}$.
In the next result, we show that even if a star-dependent matrix is not $f$-symmetric with respect to any $f$, units provide invariant subspaces.
Theorem 2.12.
Let $H$ be a hypergraph and $M_{H}$ be a star-dependent matrix indexed by $V(H)$. For any $W\in\mathfrak{U}(H)$ with $|W|\geq 2$, the vector space $C_{W}$ is an invariant subspace of $M_{H}$.
Proof.
For any $x\in C_{W}$, it is enough to show $M_{H}x\in C_{W}$. Since $x\in C_{W}$, we have $x(i)=c_{x}$ for all $i\in W$. $M_{H}=\left(m_{ij}\right)_{i,j\in V(H)}$ is a star-dependent matrix. Thus, if $i,j\in W$, then for any $k\in V(H)\setminus\{i,j\}$, we have $m_{ik}=m_{jk}$, and $m_{ij}=m_{ji}$. This leads us to
$(M_{H}x)(i)=\sum\limits_{k(\neq j)\in V(H)}m_{ik}x(k)+m_{ij}c_{x}=\sum\limits_{k(\neq i)\in V(H)}m_{jk}x(k)+m_{ji}c_{x}=(M_{H}x)(j).$
Therefore, $M_{H}x\in C_{W}$ for all $x\in C_{W}$, and this completes the proof. ∎
The above theorem indicates that units are clusters of vertices such that once they reach a consensus under the action of a star-dependent matrix, the consensus is retained in the subsequent steps.
2.2. Edge-sum Matrices and their Invariant Subspace
In 2.7, we have introduced star-dependent matrices. Now, we need to introduce a special subclass of star-dependent matrices, where the functions of stars, $f$, and $h$ are sums over the stars.
Definition 2.13 (Edge-sum matrix).
A star-dependent matrix $M_{H}=\left(m_{ij}\right)_{i,j\in V(H)}$ is called an edge-sum matrix if there exists two maps $p_{1}:E(H)\to\mathbb{R}$, and $p_{2}:E(H)\to\mathbb{R}$, such that,
$g(E_{i}(H),E_{j}(H))=\sum\limits_{e\in E_{i}(H)\cap E_{j}(H)}p_{1}(e),\text{~{}and~{}}h(E_{i}(H))=\sum\limits_{e\in E_{i}(H)}p_{2}(e),$
where $f$ and $g$ are the diagonal and non-diagonal star-functions of the star-dependent matrix $M_{H}$. For $i,j\in V(H)$, if there exists a bijection $b_{ij}:E_{i}(H)\to E_{j}(H)$ such that for any $k\notin W_{E_{i}}\cup W_{E_{j}}$
$\sum\limits_{e\in E_{i}(H)\cap E_{k}(H)}p_{1}(e)=\sum\limits_{b_{ij}(e)\in E_{j}(H)\cap E_{k}(H)}p_{1}(b_{ij}(e))\text{~{}and~{}}\sum\limits_{e\in E_{i}(H)}p_{2}(e)=\sum\limits_{b_{ij}(e)\in E_{j}(H)}p_{2}(b_{ij}(e)),$
then we say $b_{ij}$ is compatible with the edge-sum matrix $M_{H}$. We refer to $p_{1}$ and $p_{2}$ as non-diagonal and diagonal edge-weight, respectively.
Thus, if $M_{H}=\left(m_{ij}\right)_{i,j\in V(H)}$ is an edge-sum matrix and there exists a bijection $b_{ij}:E_{i}(H)\to E_{j}(H)$, compatible with $M_{H}$ then $m_{ik}=m_{jk}$, and $m_{ki}=m_{kj}$. A bijection between stars, compatible with an edge-sum matrix, leads us to the following equivalence relation for a hypergraph $H$ and an edge-sum matrix $M_{H}$.
$\mathfrak{R}_{2}(H,M_{H})=\{(i,j)\in V(H)\times V(H):\text{~{}there exists a bijection~{}}b_{ij}\text{~{}compatible with~{}}M_{H}.\}$
Now, we show that $\mathfrak{R}_{2}(H,M_{H})$ is finer than $\mathfrak{R}_{M_{H}}$.
Theorem 2.14.
Let $H$ be a hypergraph. If $M_{H}$ is an edge-sum matrix, then $\mathfrak{R}_{2}(H,M_{H})$ is finer than $\mathfrak{R}_{M_{H}}$.
Proof.
Since $M_{H}$ is an edge-sum matrix, if $(i,j)\in\mathfrak{R}_{2}(H,M_{H})$, then $m_{ik}=m_{jk}$, and $m_{ki}=m_{kj}$ for all $k\in V(H)$. Thus, proceeding like the proof of 2.10, we show $(i,j)\in\mathfrak{R}_{M_{H}}$. ∎
2.4 provides us with an invariant subspace of edge-sum matrices, which is stated in the corollary below
Corollary 2.15.
Let $H$ be a hypergraph, $f:V(H)\to(0,\infty)$ be a function, and $M_{H}$ be an $f$-symmetric, edge-sum matrix. For any $\mathfrak{R}_{2}(H,M_{H})$-equivalence class $W$ with $|W|\geq 2$, the vector space $C_{W}^{f}$ is an invariant subspace of $M_{H}$.
Note that if a star-dependent matrix $M_{H}$ is not an edge-sum matrix but satisfies $m_{ik}=m_{jk}$, and $m_{ki}=m_{kj}$ for all $i,j\in W(\subseteq V(H))$, then there is an invariant subspace of $M_{H}$ associated with $W$. We show such an instance in 3.2.
Example 2.16.
Let $W_{E}$ be a unit in a hypergraph $H$, and $i,j\in W_{E}$, then $(i,j)\in\mathfrak{R}_{2}(H,M_{H})$ for any edge-sum matrix $M_{H}$ related to $H$. Here, $b_{ij}$ is the identity map on $E_{i}(H)=E=E_{j}(H)$.
The previous example shows that the equivalence relation $\mathfrak{R}_{s}(H)$ is finer than the equivalence relation $\mathfrak{R}_{2}(H,M_{H})$ for any edge-sum matrix $M_{H}$ related to $H$. Thus, each $\mathfrak{R}_{2}(H,M_{H})$-equivalence class is a union of units. Now, if two units are in the same $\mathfrak{R}_{2}(H,M_{H})$-equivalence class, there should be a structural symmetry between the pair of units. The following notion shows a possible structural symmetry between a pair of units belonging to a $\mathfrak{R}_{2}(H,M_{H})$-equivalence class.
Twin units: Let $H$ be a hypergraph. A pair of $W_{E},W_{F}\in\mathfrak{U}(H)$, are referred to as twin units [2] if there exists a bijection $f_{EF}:E\to F$ defined by
$f_{EF}(e)=(e\setminus W_{E})\cup W_{F}.$
We refer to $f_{EF}$ as the canonical bijection of the twin units $W_{E}$, and $W_{F}$.
Example 2.17.
In the hypergraph we have considered in 2.6(see Figure 1), $W_{E_{1}}$, and $W_{E_{2}}$ are a pair of twin units.
In 2.8(2), we have considered the adjacency matrix $A^{(2)}$, which is an edge-sum matrix with non-diagonal edge-weight $p_{1}:E(H)\to\mathbb{R}$ is defined by $p_{1}(e)=\frac{1}{|e|-1}$ for all $e\in E(H)$, and the diagonal weight $p_{2}:E(H)\to\mathbb{R}$ defined by $p_{2}(e)=0$ for all $e\in E(H)$. Let $W_{E_{r}}$, and $W_{E_{s}}$ be a pair of twin units in $H$ with $|W_{E_{r}}|=|W_{E_{s}}|$. For two distinct $i,j\in W_{E_{r}}\cup W_{E_{s}}$, if both $i,j\in W_{E_{r}}$ or $i,j\in W_{E_{s}}$ then $b_{ij}$ is the identity map. If $i\in W_{E_{r}}$, and $j\in W_{E_{s}}$ then we set $b_{ij}=f_{E_{r}E_{s}}$, the canonical bijection of the pair of twin units $W_{E_{r}}$, and $W_{E_{s}}$. Thus, $(i,j)\in\mathfrak{R}_{2}(H,A^{(2)})$. Therefore, there exists an $\mathfrak{R}_{2}(H,A^{(2)})$-equivalence class $W$, such that $W_{E_{r}}\cup W_{E_{s}}\subseteq W$. For a positive-valued functions $f:V(H)\to(0,\infty)$, by 2.15, $C^{f}_{W}$ is an invariant subspace of $A^{(2)}$. Similarly, for the other matrices considered in 2.8 and 2.9, we have similar invariant subspaces. Now, the union of twin units is a set of vertices with similar stars. Now, we show the union of equipotent twin units leads us to some invariant subspace because of structural symmetry. We refer to an edge-sum matrix $M_{H}$ as edge-cardinality matrix if the weights $p_{1}$ and $p_{2}$ only depend on the cardinality of the hyperedges. That is, there exists two functions $q_{1}:\mathbb{N}\to\mathbb{R}$, and $q_{2}:\mathbb{N}\to\mathbb{R}$ such that $p_{1}(e)=q_{1}(|e|)$, and $p_{2}(e)=q_{2}(|e|)$ for all $e\in E(H)$.
Theorem 2.18.
Let $H$ be a hypergraph, and $M_{H}=(m_{ij})_{i,j\in V(H)}$ be an edge-cardinality matrix associated with $H$. If $W_{E}$, and $W_{F}$ are a pair of distinct twin units such that $|W_{E}|=|W_{F}|$, then there exists an eigenvalue $\lambda$ of $M_{H}$ with an eigenvector $\chi_{W_{E}}-\chi_{W_{F}}$.
Proof.
For any $r\in W_{E}$,
$\sum\limits_{k\in W_{E}}m_{rk}=(|W_{E}|-1)\sum\limits_{e\in E}q_{1}(|e|)+\sum\limits_{e\in E}q_{2}(|e|)=\mu_{W_{E}}\text{~{}say~{}}.$
Similarly, we have the constant $\mu_{W_{F}}$ associated with $W_{F}$. If two units $W_{E}$, and $W_{F}$ are distinct, that is $W_{E}\neq W_{F}$, then either $W_{E}\setminus W_{F}\neq\emptyset$ or $W_{F}\setminus W_{E}\neq\emptyset$. Without loss of generality let $W_{E}\setminus W_{F}\neq\emptyset$, and $i\in W_{E}\setminus W_{F}$. Thus, $E=E_{i}(H)\neq F$. Therefore, $W_{E}\neq W_{F}$ implies $W_{E}\cap W_{F}=\emptyset$, otherwise if $W_{E}\cap W_{F}\neq\emptyset$, and $k\in W_{E}\cap W_{F}$, then $E=E_{k}(H)=F$, a contradiction. Given that $|W_{E}|=|W_{F}|$, the two units are either identical or have no common elements, and $\mathfrak{f}_{EF}(e)=(e\setminus W_{E})\cup W_{F}$, it follows that $|e|=|\mathfrak{f}_{EF}(e)|$. Thus,
$\mu_{W_{E}}$ $=(|W_{E}|-1)\sum\limits_{e\in E}q_{1}(|e|)+\sum\limits_{e\in E}q_{2}(|e|)$
$=(|W_{F}|-1)\sum\limits_{e\in E}q_{1}(|\mathfrak{f}_{EF}(e)|)+\sum\limits_{e\in E}q_{2}(|\mathfrak{f}_{EF}(e)|)=\mu_{W_{F}}.$
Suppose that $y_{EF}=\chi_{W_{E}}-\chi_{W_{F}}$. We consider the following cases to compute $(M_{H}y_{EF})$.
  1. (1)
    For $r\in W_{E}$, if $k\in W_{F}$, then $E_{r}(H)\cap E_{k}(H)=\emptyset$, and thus, $m_{rk}=\sum\limits_{E_{r}(H)\cap E_{k}(H)}p_{1}(e)=0$. Since $y_{EF}(k)=0$ for all $k\in V(H)\setminus(W_{E}\cup W_{F})$
    $(M_{H}y_{EF})(r)=\sum\limits_{k\in W_{E}}m_{rk}=\mu_{W_{E}}$
  2. (2)
    Similarly, for $r\in W_{F}$, $(M_{H}y_{EF})(r)=-\mu_{W_{F}}$.
  3. (3)
    Since $|W_{E}|=|W_{F}|$, If $r\in V(H)\setminus(W_{E}\cup W_{F})$, then for any $k\in W_{E},s\in W_{F}$,
    $m_{rk}=\sum\limits_{e\in E_{r}(H)\cap E}q_{1}(|e|)=\sum\limits_{\mathfrak{f}_{EF}(e)\in E_{r}(H)\cap F}q_{1}(|\mathfrak{f}_{EF}(e)|)=m_{rs}.$
    Therefore, for $r\in V(H)\setminus(W_{E}\cup W_{F})$,
    $(M_{H}y_{EF})(r)=\sum\limits_{k\in W_{E}}m_{rk}-\sum\limits_{s\in W_{F}}m_{rs}=0.$
Thus,
$(M_{H}y_{EF})(r)=\begin{cases}\phantom{-}\mu_{W_{E}}&\text{~{}if~{}}r\in W_{E}\\ -\mu_{W_{F}}&\text{~{}if~{}}r\in W_{F}\\ \phantom{-}0&\text{~{}otherwise.}\end{cases}$ (2.3)
Since $\mu_{W_{E}}=\mu_{W_{F}}$, the result follows. ∎
Now, we show that the union of a pair of equipotent twin units provides us with an invariant subspace of any edge-cardinality matrix associated with a hypergraph.
Theorem 2.19.
Let $H$ be a hypergraph, $f:V(H)\to(0,\infty)$ be a function, and $M_{H}$ be an $f$-symmetric, edge-cardinality matrix. If $W_{E}$, and $W_{F}$ are a pair of distinct twin units with $|W_{E}|=|W_{F}|$, then $C^{f}_{W}$ is an invariant subspace of $M_{H}$, where $W=W_{E}\cup W_{F}$.
Proof.
Since each edge cardinality matrix is a star dependent matrix, by 2.10, and 2.2, there exists $\lambda_{W_{E}},\lambda_{W_{F}}\in\sigma_{M_{H}}$ such that, $S_{W_{E}}$ and $S_{W_{F}}$ are subspaces of $V_{\lambda_{W_{E}}}(M_{H})$, and $V_{\lambda_{W_{F}}}(M_{H})$ respectively. Thus, both $S_{W_{E}}$, and $S_{W_{F}}$ are invariant subspaces of $M_{H}$. By 2.18, the vector space $\langle\chi_{W_{E}}-\chi_{W_{F}}\rangle$, spanned by the vector $\chi_{W_{E}}-\chi_{W_{F}}$ is also an invariant subspace of $M_{H}$. $M_{H}$ is $f$-symmetric. Thus, $C^{f}_{W}=(S_{W_{E}}\oplus S_{W_{F}}\oplus\langle\chi_{W_{E}}-\chi_{W_{F}}\rangle)^{(\perp,f)}$ is an invariant subspace of $M_{H}$, where $W=W_{E}\cup W_{F}$. ∎
We have seen that a pair of twin units is a pair of units in the same $\mathfrak{R}_{2}(H,M_{H})$-equivalence class for some specific matrices $M_{H}$ related to $H$. Now, we show an $\mathfrak{R}_{2}(H,M_{H})$-equivalence class is not necessarily a union of twin units, and there may be other structural symmetries as well.
{tikzpicture}
Figure 2. A hypergraph $H$ with its vertex set partitioned in units
Example 2.20.
Consider the hypergraph $H$ in Figure 2 with $V(H)=\{1,2,\ldots,16\}$, and $E(H)=\{e,f,g,h,i\}$. Where $e=\{1,2,3,4,11,16\}$, $f=\{3,4,5,6,12,15\}$, $g=\{5,6,7,8,10,14\}$, $h=\{1,2,7,8,13\}$, $i=\{2,3,5,7\}$ are the hyperedges in $H$. Here $\mathfrak{U}(H)=\{W_{E_{i}}:i=1,2,\ldots,12\}$ with $W_{E_{1}}=\{2\}$, $W_{E_{2}}=\{3\}$, $W_{E_{3}}=\{5\}$, $W_{E_{4}}=\{7\}$, $W_{E_{5}}=\{11,16\}$, $W_{E_{6}}=\{12,15\}$, $W_{E_{7}}=\{10,14\}$, $W_{E_{8}}=\{9,13\}$, $W_{E_{9}}=\{1\}$, $W_{E_{10}}=\{4\}$, $W_{E_{11}}=\{6\}$, $W_{E_{12}}=\{8\}$. Here, no two of the four units, $W_{E_{1}},W_{E_{2}},W_{E_{3}},W_{E_{4}}$ form a pair of twin units. Consider $W=W_{E_{1}}\cup W_{E_{2}}\cup W_{E_{3}}\cup W_{E_{4}}$. Now, we define all possible $b_{ij}$ for $i,j(\neq i)\in W$. Consider $b_{23}:E_{2}(H)\to E_{3}(H)$ as $e\mapsto e$, $i\mapsto i$, and $h\mapsto f$; $b_{35}:E_{3}(H)\to E_{5}(H)$ as $e\mapsto g$, $i\mapsto i$, and $f\mapsto f$; $b_{57}:E_{5}(H)\to E_{7}(H)$ as $g\mapsto g$, $i\mapsto i$, and $f\mapsto h$; $b_{72}:E_{7}(H)\to E_{2}(H)$ as $g\mapsto e$,$i\mapsto i$, and $h\mapsto h$; $b_{37}:E_{3}(H)\to E_{7}(H)$ as $e\mapsto g$,$i\mapsto i$, and $f\mapsto h$; $b_{25}:E_{2}(H)\to E_{5}(H)$ as $e\mapsto g$,$i\mapsto i$, and $h\mapsto f$. Take $b_{ji}=b_{ij}^{-1}$. Let $M_{H}$ be the matrix $A^{(2)}$ associated with $H$. For $p\neq q$, the $(p,q)$-th element of $A^{(2)}$ is $\sum\limits_{e\in E_{p}(H)\cap E_{q}(H)}\frac{1}{|e|-1}$. The non-diagonal weight $p_{1}(e)=\frac{1}{|e|-1}$, a function of hyperedge cardinality. These bijections $\{b_{ij}:{i,j(\neq i)\in W}\}$, being cardinality preserving, are compatible with $A^{(2)}$ associated with $H$. Thus, $W$ is an $\mathfrak{R}_{2}(H,A^{(2)})$-equivalence class. Since $A^{(2)}$ is symmetric with respect to the usual inner product, by 2.15, $C_{W}$ is an invariant subspace of $A^{(2)}$.
Similarly, $W^{\prime}=\{1,3,4,6\}$ is an $\mathfrak{R}_{2}(H,A^{(2)})$-equivalence class, which is not a union of twin units.
3. Application
In the previous section, we have seen the symmetry represented by equivalence relations on the set of vertices provides us with invariant subspaces of specific matrices associated with hypergraphs. In this section, we demonstrate that the consistent behaviours observed in specific clusters of vertices during random walks and diffusion on hypergraphs can be attributed to invariant subspaces.
3.1. Random walk on hypergraphs
[…]
$x_{t+1}(u)=\sum\limits_{v\in V(H)}x_{t}(v)P_{vu},\text{~{}that is,~{}}x_{t+1}=P^{T}x_{t}.$
Here, we consider only non-lazy random walk, that is, for any $u$, if $r(t)=u$ for some $t\in\mathbb{Z}_{+}$, then $r(t+1)\neq u$. In [1], it has been shown that $P=\mathcal{A}$, the normalized adjacency matrix, described in 2.8(3). Now we use 2.12 to prove that if at a particular time $s$, the values of probability distribution $x_{s}$ become equal in all the vertices in a unit, then equality will be retained in the subsequent time steps.
Theorem 3.1.
Let $H$ be a hypergraph, and $r:\mathbb{Z}_{+}\to V(H)$ be a random walk on $H$ with probability distribution $x_{t}$ at any time $t\in\mathbb{Z}_{+}$. For any unit $W_{E}\in\mathfrak{U}(H)$, if $x_{s}(i)=p_{s}$, a constant for all $i\in W_{E}$ for some $s\in\mathbb{Z}_{+}$, then $x_{t}(i)=p_{t}$, a constant for all $i\in W_{E}$ for all $t\geq s$.
Proof.
If $x_{s}(i)=p_{s}$, a constant for all $i\in W$ for some $s\in\mathbb{Z}_{+}$, then $x_{s}\in C_{W_{E}}$. Since $\mathcal{A}$, the normalized adjacency matrix, described in 2.8(3) is a star dependent matrix (see 2.8), by 2.12, $x_{t}\in C_{W_{E}}$ for all $t\geq s$. Thus, the theorem follows. ∎
Now, we show that a similar result is also true for a union of equipotent twin units.
Theorem 3.2.
Let $H$ be a hypergraph, and $r:\mathbb{Z}_{+}\to V(H)$ be a random walk on $H$ with probability distribution $x_{t}$ at any time $t\in\mathbb{Z}_{+}$. For any pair of twin units $W_{E},W_{F}\in\mathfrak{U}(H)$ with the canonical bijection ${f}_{EF}:E\to F$, if $|W_{E}|=|W_{F}|$ and $x_{s}(r)=p_{s}$ for all $r\in W_{E}\cup W_{F}$ for some $s\in\mathbb{Z}_{+}$, then $x_{t}(r)=p_{t}$ for all $r\in W_{E}\cup W_{F}$ for all $t\geq r$.
Proof.
Let $W_{E},W_{F}\in\mathfrak{U}(H)$ be a pair of twin units with $|W_{E}|=|W_{F}|$ with the canonical bijection ${f}_{EF}:E\to F$. Since $|W_{E}|=|W_{F}|$, we have $|{f}_{EF}(e)|=|e|$ for all $e\in E$. For any $i\in W_{E}$, and $j\in W_{F}$, we have $\mathcal{A}_{ki}=\frac{1}{|E_{k}(H)|}\sum\limits_{e\in E_{k}(H)\cap E_{i}(H)}\frac{1}{|e|-1}=\frac{1}{|E_{k}(H)|}\sum\limits_{\mathfrak{f}_{EF}(e)\in E_{k}(H)\cap E_{j}(H)}\frac{1}{|e|-1}=\mathcal{A}_{kj}$, where $k\in V(H)\setminus(W_{E}\cup W_{F})$. For any $i,k(\neq i)\in W_{E}$, and $j,k^{\prime}(\neq j)\in W_{F}$ we have $\mathcal{A}_{ki}=\frac{1}{|E|}\sum\limits_{e\in E}\frac{1}{|e|-1}=\frac{1}{|F|}\sum\limits_{\mathfrak{f}_{EF}(e)\in F}\frac{1}{|e|-1}=\mathcal{A}_{k^{\prime}j}$. Thus, if $x_{s}(r)=p_{s}$ for all $r\in W_{E}\cup W_{F}$ for some $s\in\mathbb{Z}_{+}$, then for any pair $i\in W_{E}$, and $j\in W_{F}$ we have
$x_{s+1}(i)$ $=\sum\limits_{k\in V(H)}x_{s}(k)\mathcal{A}_{ki}$
$=\sum\limits_{k(\neq i)\in W_{E}}p_{s}\mathcal{A}_{ki}+\sum\limits_{k\in V(H)\setminus(W_{E}\cup W_{F})}x_{s}(k)\mathcal{A}_{ki}$
$=\sum\limits_{k^{\prime}(\neq j)\in W_{E}}p_{s}\mathcal{A}_{k^{\prime}j}+\sum\limits_{k\in V(H)\setminus(W_{E}\cup W_{F})}x_{s}(k)\mathcal{A}_{kj}=x_{s+1}(j).$
If both $i,j$ are either in $W_{E}$ or belongs to $W_{F}$, then, since $x_{s}$ is constant on both the unit, by 3.1, $x_{t}$ would be constant for all $t\geq s$. Now, when $i\in W_{E}$, and $j\in W_{F}$, we have $x_{s+1}(i)=x_{s+1}(j)$. Therefore, $x_{t}$ is constant on $W_{E}\cup W_{F}$ for all $t\geq s$. ∎
3.2. Diffusion on hypergraphs
We have described some Laplacians related to a hypergraph $H$ in 2.9. Let $L_{H}$ be any Laplacian described in that example. If the hypergraph is connected, then the only non-positive eigenvalue of $L_{H}$ is 00 with multiplicity $1$. The eigenspace $V_{0}(L_{H})$ corresponding to the 00 eigenvalue is the $1$-dimensional subspace generated by $\chi_{V(H)}$, the characteristic function of $V(H)$. Thus, any solution of the differential equation (˙$\dot{(}x)=L_{H}x$ tends to $c\chi_{V(H)}$ for some real constant $c$. For this reason, the Laplacian is said to have a diffusive influence on the hypergraph $H$. The following difference equation can describe the corresponding diffusion process.
$x_{t+1}=Lx_{t}.$ (3.1)
It has been observed that if the initial state $x_{0}$ is of the form $c\chi_{V(H)}+\delta y$ for sufficiently small $\delta>0$, and $y\in\mathbb{R}^{V(H)}$, we have $\lim\limits_{t\to\infty}x_{t}=c_{0}\chi_{V(H)}$, that is, the states of all the nodes will reach to a consensus under the influence of $L_{H}$. Before reaching a complete consensus, we observed some clusters of local consensus. While exploring the reason for the small clusters of consensus, we find units preserve consensus.
Theorem 3.3.
Let $H$ be a hypergraph, and $W\in\mathfrak{U}(H)$. If for some $s\in\mathbb{Z}_{+}$, the $s$-th stage of the trajectories governed by Equation 3.1, $x_{s}$ is such that $x_{s}(i)=c_{s}$, a constant for all $i\in W$, then $x_{t}(i)=c_{t}$, a constant for all $i\in W$ for all $t\geq s$.
Proof.
Since $L_{H}$ is a star-dependent matrix, the result follows from 2.12. ∎
A similar result can be proved for a union of equipotent twin units. The proof of the result is similar to the proof of 3.2; thus, we omit the proof.
Theorem 3.4.
Let $H$ be a hypergraph, and $W_{E},W_{F}\in\mathfrak{U}(H)$ with the canonical bijection $\mathfrak{f}_{EF}:E\to F$. If $|W_{E}|=|W_{F}|$, and for some $s\in\mathbb{Z}_{+}$, the $s$-th stage of the trajectories governed by Equation 3.1, $x_{s}$ is such that $x_{s}(i)=c_{s}$, a constant for all $i\in W_{E}\cup W_{F}$, then $x_{t}(i)=c_{t}$, a constant for all $i\in W_{E}\cup W_{F}$, for all $t\geq s$.
Acknowledgement
The author, Samiron Parui, expresses gratitude to the National Institute of Science Education and Research for their financial support through the Department of Atomic Energy plan project RIA 4001 (NISER).